.A box contains four blue and three red marbles. Michelle picks



four marbles at random from this box. If Z is the random variable



representing the number of blue marbles picked from the box,



construct the probability distribution of the random variable Z

1
Expert's answer
2022-02-03T18:11:14-0500

Solution:

four blue and three red marbles

Total = 4+3 = 7

Z = the number of blue marbles picked

Z = {0,1,2,3,4}

"p=\frac 17, q=\frac 67, n=4" (p is the success getting a blue marble).

"P(Z=0)=^4C_0(\frac 17)^0(\frac 67)^4=\frac {1296}{2401}
\\ P(Z=1)=^4C_1(\frac 17)^1(\frac 67)^3=\frac {864}{2401}
\\ P(Z=2)=^4C_2(\frac 17)^2(\frac 67)^2=\frac {216}{2401}
\\P(Z=3)=^4C_3(\frac 17)^3(\frac 67)^1=\frac {24}{2401}
\\ P(Z=4)=^4C_4(\frac 17)^4(\frac 67)^0=\frac {1}{2401}"

Thus, the probability distribution of Z is:

Z P(Z)

0 1296/2401

1 864/2401

2 216/2401

3 24/2401

4 1/2401